OpenStudy (anonymous):

use mathematical induction 2^n>n^3

myininaya (myininaya):

counterexample: n=2 2^2=4 2^3=8 4 is not greater than 8

OpenStudy (anonymous):

probably missing that a condition that n>= 10

OpenStudy (anonymous):

oh:: 2^n>n^3 for every integer n>=10

OpenStudy (zarkon):

\[(n+1)^3=n^3+3n^2+3n+1\] \[3n^2+3n+1\leq 3n^2+3n^2+n^2=7n^2\leq n^3\]

OpenStudy (anonymous):

wat is leq

myininaya (myininaya):

i would have done: \[n=10: 2^{10}>10^3\] now assume for some k>10 the following holds: \[n=k: 2^k>k^3\] now we want to show it is true k+1 \[n=k+1: 2^{k+1}=2^k2=2\cdot2^k>2 \cdot k^3>k^3\]

OpenStudy (zarkon):

don't you need \[2^{k+1}>(k+1)^3\]